September 9, 2003

Dear Fellow Math Educator:

We take pleasure in inviting you and your fellow math colleagues and students to join the NKU math community for dinner and an after-dinner talk with Ed Burger, Mathematics Teacher, Writer, and Popularizer, and winner of major Mathematical Association of America Teaching Awards.

On Monday, October 13th, Dr. Burger will be speaking at NKU on the topic “Magic With Mathematics” (see attached flyer). The occasion is the third annual Sehnert lecture. You may have attended previous years’ lectures, by Underwood Dudley and Ivars Peterson; this year’s lecture promises to be equally entertaining. Our speaker has indicated that this talk will be accessible (and even fun!) for high school students, so we encourage you to bring your students to what promises to be a wonderful talk.

In addition, we would like you and some of your students to join us for a reception and buffet dinner preceding the talk, to begin at 5:30, where you can meet Dr. Burger, NKU Math and Computer Science faculty, current NKU students, and other high school faculty and students from the area. Dinner will be served at 6:00 in the NKU ballroom (University Center) with the lecture to follow at 7:30 in the Otto Budig theater.

To reserve your place at dinner, please contact the department at 859-572-5377. If you wish to bring more than two students for the dinner, please be sure to call and check on availability of space.

If you have any questions, feel free to call me at 572-5794, or email me.

Please join us on the 13th! We believe that you will experience an enjoyable and rewarding evening.

Sincerely,

Andy Long, longa@nku.edu, Speakers Committee, NKU Department of Math and Computer Science